### v2

# Second-Order Tensorial Partial Differential Equations on Graphs

Abstract
Processing data on multiple interacting graphs is crucial for many applications, but existing approaches rely mostly on discrete filtering or first-order continuous models that dampen high frequencies and propagate information slowly. We introduce second-order tensorial partial differential equations on graphs (So-TPDEGs) and propose the first theoretically grounded framework for second-order continuous product graph neural networks. Our method exploits the separability of cosine kernels in Cartesian product graphs to enable efficient spectral decomposition while preserving high-frequency signals. We further provide rigorous analyses of stability under graph perturbations and over-smoothing, establishing a solid theoretical foundation for continuous graph learning.
Index Terms—  Partial differential equations on graphs, graph neural networks, over-smoothing, stability
1 Introduction
Tensors [1], which generalize matrices to higher dimensions, arise in diverse domains such as hyperspectral imaging [2], recommender systems [3], and spatiotemporal modeling [4]. When tensors are linked to multiple interdependent graphs, they give rise to multidomain graph data [5, 6, 7]. Unlike conventional methods in graph signal processing (GSP) [8] and graph machine learning (GML), typically considering a single graph [6], this setting requires jointly processing across several interacting graph domains. Although such data structures are increasingly common in real-world applications, the methods for learning from multidomain graph data remain scarce [9, 6, 10, 11]. Thus, developing graph learning techniques for tensorial data holds great promise for advancing the theory and applications in GSP and GML.
A common strategy for learning from multidomain graph data is to use product graphs (PGs) [9, 12, 13], a concept rooted in GSP [8], where discrete filtering operations combine information across domains. For example, the graph-time convolutional neural network (GTCNN) [9] introduces a discrete filtering mechanism to jointly process spatial and temporal information, but suffers from three key limitations: i) it requires costly grid searches to tune filter orders, ii) polynomial filters restrict receptive fields and limit long-range modeling [14, 15], and iii) it is inherently constrained to two-factor graphs (space and time). Besides, these methods inherit well-known limitations of conventional graph neural networks (GNNs) [16, 17], such as over-smoothing and over-squashing, which restrict receptive fields and long-range modeling [14].
In another line of research, continuous GNNs (CGNNs) have emerged as an effective approach to mitigate the challenges of over-smoothing and over-squashing [15]. CGNNs formulate graph learning as solving partial differential equations (PDEs) on graphs—such as the heat or wave equations—through neural architectures that approximate these solutions [15, 18]. While these frameworks have shown promise, they are mostly designed for single-graph settings. A notable exception is CITRUS [10], which extends CGNNs to multidomain data by introducing TPDEGs and using PGs. Despite its contributions, CITRUS is limited to first-order derivatives [10], which dampen high-frequency signals and slow information propagation—issues that are particularly problematic in oscillatory (periodic or high-frequency fine-grained) tasks such as spatiotemporal forecasting, heterophilic networks, and continuum mechanics simulations [19, 20, 21].
To overcome these limitations, we propose the first principled second-order TPDEGs (So-TPDEGs), which provide a stronger foundation for modeling multidomain data across multiple interacting graphs. Building on this formulation, we introduce a continuous graph learning model that efficiently solves So-TPDEGs using separable cosine oscillatory kernels on Cartesian PGs, which preserves high-frequency components. This design also allows receptive fields to be learned adaptively during training, eliminating the costly grid searches required by discrete methods. Our design avoids the computational burden of full spectral decompositions by using a small subset of eigenvalue decompositions from the factor graphs. We further provide rigorous theoretical analyses, showing that our model preserves stability under graph perturbations and controls the over-smoothing rate. Importantly, the number of learnable parameters is independent of the number of factor graphs, ensuring scalability to large and complex multidomain settings.
The key contributions of this work are summarized as follows: i) We propose So-TPDEGs as a unified framework for modeling multidomain graph data. ii) We derive a continuous graph filtering mechanism over Cartesian product spaces as a natural solution to So-TPDEGs. iii) We provide comprehensive theoretical analyses of our model, establishing its stability and demonstrating its ability to alleviate over-smoothing.
2 Methodology
Preliminaries and notation. An undirected, weighted graph $\mathcal{G}$ with $N$ vertices is represented as $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{A}\}$, where $\mathcal{V}$ and $\mathcal{E}$ denote the sets of vertices and edges, respectively. The adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ characterizes the graph structure, with $\mathbf{A}_{ij}=\mathbf{A}_{ji}\geq 0$ indicating the connection strength between vertices $i$ and $j$. We assume $\mathbf{A}_{ii}=0$ for all $i$, meaning the graph contains no self-loops. The graph Laplacian $\mathbf{L}\in\mathbb{R}^{N\times N}$ is defined as $\mathbf{L}=\mathbf{D}-\mathbf{A},$ where $\mathbf{D}=\mathrm{diag}(\mathbf{A}\mathbf{1})$ is the diagonal degree matrix and $\mathbf{1}$ is the all-ones vector. A graph signal is defined as a mapping $x:\mathcal{V}\to\mathbb{R}$, which assigns a real value to each node. Such a signal can be represented in vector form as $\mathbf{x}=[x_{1},\ldots,x_{N}]^{\top}$. For a general vector $\mathbf{a}=[a_{1},\ldots,a_{N}]^{\top}\in\mathbb{R}^{N\times 1}$, we define the element-wise cosine as $\cos(\mathbf{a})=[\cos(a_{1}),\ldots,\cos(a_{N})]^{\top}.$ We denote the vectorization operator by $\mathrm{vec}(\cdot)$. Finally, using the Kronecker product $\otimes$, the Kronecker sum [22] (also referred to as the Cartesian product) $\oplus$, and the Laplacian factors $\{\mathbf{L}_{p}\in\mathbb{R}^{N_{p}\times N_{p}}\}_{p=1}^{P}$, we introduce the following definitions:
$\begin{split}&\downarrow\oplus_{p=1}^{P}{\mathbf{L}_{p}}\vcentcolon=\mathbf{L}_{P}\oplus\ldots\oplus\mathbf{L}_{1},\\ &\otimes_{p=1}^{P}{\mathbf{L}_{p}}\vcentcolon=\mathbf{L}_{1}\otimes\ldots\otimes\mathbf{L}_{P},\\ &\downarrow\otimes_{p=1}^{P}{\mathbf{L}_{p}}\vcentcolon=\mathbf{L}_{P}\otimes\ldots\otimes\mathbf{L}_{1}.\end{split}$ (1)
Here, $\mathbf{I}_{n}$ denotes the identity matrix of size $n$. An illustration of a Cartesian graph product between three factor graphs is shown in Figure 1(a). We define a $D$-dimensional tensor as $\underline{\mathbf{U}}\in\mathbb{R}^{N_{1}\times\cdots\times N_{D}}$. The matrix $\underline{\mathbf{U}}_{(i)}\in\mathbb{R}^{N_{i}\times\left[\prod_{r=1,r\neq i}^{D}N_{r}\right]}$ represents the $i$-th mode matricization (unfolding) of the tensor $\underline{\mathbf{U}}$ [1]. The mode-$i$ tensor multiplication of a tensor $\underline{\mathbf{U}}$ by a matrix $\mathbf{X}\in\mathbb{R}^{m\times N_{i}}$ is defined as $\underline{\mathbf{G}}=\underline{\mathbf{U}}\times_{i}\mathbf{X},$ where $\underline{\mathbf{G}}\in\mathbb{R}^{N_{1}\times\cdots\times N_{i-1}\times m\times N_{i+1}\times\cdots\times N_{D}}$ [1]. All the proofs of theorems and propositions of this paper are provided in the Appendix.
2.1 Second-Order Continuous Product Graph Neural Networks
We first propose the second-order generative PDE for PGs as follows:
Definition 2.1 (So-TPDEG).
Let $\tilde{\underline{\mathbf{U}}}_{t}\in\mathbb{R}^{N_{1}\times N_{2}\times\cdots\times N_{P}}$ denote a multidomain tensor whose elements vary as functions of time $t$. We further define the So-TPDEG with $\{\mathbf{L}_{p}\in\mathbb{R}^{N_{p}\times N_{p}}\}_{p=1}^{P}$ as the set of Laplacians corresponding to each factor graph, as follows:
$\frac{\partial^{2}\underline{\mathbf{U}}_{t}}{\partial\>t^{2}}=-\sum_{i=1}^{P}\underline{\mathbf{U}}_{t}\times_{i}\mathbf{L}_{i}^{2}-2\sum_{1\leq i<j\leq P}\underline{\mathbf{U}}_{t}\times_{i}\mathbf{L}_{i}\times_{j}\mathbf{L}_{j}.$ (2)
Here, there are two important key differences with the first-order TPDEGs [10]: (i) The first term in the RHS applies $\mathbf{L}^{2}_{i}$, compared to TPDEG, which exploits $\mathbf{L}_{i}$. Therefore, So-TPDEG potentially searches for further neighborhoods. (ii) The second term in the RHS of (2) models the cross-interaction between factor graphs, compared to the TPDEG relying only on the independencies of the domain-specific factor graphs.
Fig. 1: (a) Cartesian product of three factor graphs. (b) Continuous product graph function in So-TPDEG on the tensor $\underline{\mathbf{U}}$.
Theorem 2.2.
Assuming $\tilde{\underline{\mathbf{U}}}_{0}$ as the initial value of $\tilde{\underline{\mathbf{U}}}_{t}$, the solution to the So-TPDEG in (2) (with the assumption of $\left.\frac{\partial\underline{\mathbf{U}_{t}}}{\partial t}\right|_{t=0}=0$) is given by:
$\begin{split}&\tilde{\underline{\mathbf{U}}}_{t}=\sum_{\begin{subarray}{c}\boldsymbol{\epsilon}\in\{0,1\}^{P}\\ |\boldsymbol{\epsilon}|\,\text{even}\end{subarray}}(-1)^{|\boldsymbol{\epsilon}|/2}\;\tilde{\underline{\mathbf{U}}}_{0}\times_{1}g_{t}^{(\epsilon_{1})}(\mathbf{L}_{1})\cdots\times_{P}g_{t}^{(\epsilon_{P})}(\mathbf{L}_{P}),\end{split}$ (3)
where $g_{t}^{(0)}(\mathbf{L}_{i})=\cos(t\mathbf{L}_{i}),\>\>g_{t}^{(1)}(\mathbf{L}_{i})=\sin(t\mathbf{L}_{i})$, $\boldsymbol{\epsilon}=(\epsilon_{1},\dots,\epsilon_{P})\in\{0,1\}^{P}$, with $\{0,1\}^{P}$ being the Cartesian product between $P$ sets of $\{0,1\}$. Besides, $|\boldsymbol{\epsilon}|:=\sum_{i=1}^{P}\epsilon_{i}$.
Let $\underline{\mathbf{U}}_{l}\in\mathbb{R}^{N_{1}\times\ldots\times N_{P}\times F_{l}}$ be the input tensor, $\mathbf{W}_{l}\in\mathbb{R}^{F_{l}\times F_{l+1}}$ a matrix of learnable parameters, $t_{l}$ the learnable graph receptive field, and $f(\underline{\mathbf{U}}_{l})\in\mathbb{R}^{N_{1}\times\ldots\times N_{P}\times F_{l+1}}$ the output tensor. Using Theorem 2.2, we can define the core function of our framework as follows:
$\begin{split}&f(\underline{\mathbf{U}}_{l})=\\ &\sum_{\begin{subarray}{c}\boldsymbol{\epsilon}\in\{0,1\}^{P}\\ |\boldsymbol{\epsilon}|\,\text{even}\end{subarray}}(-1)^{|\boldsymbol{\epsilon}|/2}\;\underline{\mathbf{U}}_{l}\times_{1}g_{t_{l}}^{(\epsilon_{1})}(\mathbf{L}_{1})\cdots\times_{P}g_{t_{l}}^{(\epsilon_{P})}(\mathbf{L}_{P})\times_{P+1}\mathbf{W}^{\top}_{l}.\end{split}$ (4)
Thus, the core graph filtering operation is obtained on a PG as follows (see Figure 1(b) for a high-level picture):
Proposition 2.3.
The core function of our framework in (4) can be rewritten as:
$\left[f(\underline{\mathbf{U}}_{l})_{(P+1)}\right]^{\top}=\cos(t_{l}\mathbf{L}_{\diamond}){[\underline{\mathbf{U}}_{l}}_{(P+1)}]^{\top}\mathbf{W}_{l},$ (5)
where $\mathbf{L}_{\diamond}\vcentcolon=\downarrow\oplus_{p=1}^{P}{\mathbf{L}_{p}}$ is the Laplacian of the Cartesian PG.
Proposition 2.3 provides the theoretical basis for the implementation of our framework. To compute the solution efficiently, we can use the spectral decompositions of the factor and product Laplacians $\{\mathbf{L}_{p}=\mathbf{V}_{p}\boldsymbol{\Lambda}_{p}\mathbf{V}_{p}^{\top}\}_{p=1}^{P}$ and $\{\mathbf{L}_{\diamond}=\mathbf{V}_{\diamond}\boldsymbol{\Lambda}_{\diamond}\mathbf{V}_{\diamond}^{\top}\}$, respectively. Here, the eigenvectors of the PG are obtained as $\mathbf{V}_{\diamond}=\downarrow\otimes_{p=1}^{P}\mathbf{V}_{p}$, and the corresponding eigenvalues follow $\boldsymbol{\Lambda}_{\diamond}=\downarrow\oplus_{p=1}^{P}\boldsymbol{\Lambda}_{p}$ [22]. For each factor graph, we choose the top $K_{p}\leq N_{p}$ eigenvalue–eigenvector pairs to reduce cost. More precisely, it can be shown [14] that Eqn. (5) can be equivalently expressed in the following form:
$\begin{split}&\left[f(\underline{\mathbf{U}}_{l})_{(P+1)}\right]^{\top}=\\ &\mathbf{V}^{(K_{p})}_{\diamond}\left(\underbrace{\overbrace{[\tilde{\boldsymbol{\lambda}}_{l}|\ldots|\tilde{\boldsymbol{\lambda}}_{l}]}^{F_{l}\>\text{times}}}_{\tilde{\boldsymbol{\Lambda}}_{l}}\odot\left({\mathbf{V}^{(K_{p})}_{\diamond}}^{\top}\left[{\underline{\mathbf{U}}_{l}}_{(P+1)}\right]^{\top}\right)\right)\mathbf{W}_{l},\end{split}$ (6)
[…]
Here, $\boldsymbol{\lambda}^{(K_{p})}_{p}\in\mathbb{R}^{K_{p}\times 1}$ and $\mathbf{V}^{(K_{p})}_{p}\in\mathbb{R}^{N_{p}\times K_{p}}$ represent the top $K_{p}\leq N_{p}$ eigenvalues and their corresponding eigenvectors of $\mathbf{L}_{p}$, selected according to the largest magnitudes of the eigenvalues. The symbol $\odot$ denotes the element-wise (Hadamard) product. Then, the output of the $(l)$-th layer in our model is expressed as ${\underline{\mathbf{U}}_{l+1}}_{(P+1)}=\sigma\left(f(\underline{\mathbf{U}}_{l})_{(P+1)}\right)$, where $\sigma(\cdot)$ denotes a suitable nonlinear activation function.
Remark 2.4.
The eigenvalue decomposition (EVD) of the PG Laplacian $\mathbf{L}_{\diamond}$ in (5) has a computational complexity of $\mathcal{O}([\prod_{p=1}^{P}N_{p}]^{3})$ in the general case. However, leveraging the separable structure of PGs significantly reduces the complexity to $\mathcal{O}(\sum_{p=1}^{P}N_{p}^{3})$ in (6), since EVD is performed independently on each factor graph. Furthermore, if we retain only the top $K_{p}$ most significant eigenvalue-eigenvector pairs [14] per factor graph, the cost of spectral decomposition is further reduced to $\mathcal{O}(N_{p}^{2}K_{p})$ for each factor, resulting in an overall complexity of $\mathcal{O}(\sum_{p=1}^{P}N_{p}^{2}K_{p})$.
2.2 Stability Analysis
[…]
Proposition 2.5.
Let $\mathbf{L}_{\diamond}$ and $\tilde{\mathbf{L}}_{\diamond}$ denote the original and perturbed Laplacian matrices of the Cartesian PG, where the perturbation for each factor graph follows the model described in (8). Under this formulation, the perturbation matrix $\mathbf{E}$ admits the Cartesian structure, such that $\mathbf{E}=\oplus_{p=1}^{P}\mathbf{E}_{p}$, and the following relationship holds:
[…]
Let $\varphi(u,t)$ and $\tilde{\varphi}(u,t)$ denote the outputs of our model under the original and perturbed Laplacian, respectively. The Laplacian $\mathbf{L}_{\diamond}$ governs the evolution of ${\underline{\mathbf{U}}_{t}}_{(P+1)}$ through the flow PDE $\frac{\partial^{2}\varphi(u,t)}{\partial t^{2}}=-\mathbf{L}_{\diamond}^{2}\varphi(u,t)$ as described in (5), with the corresponding solution $\varphi(u,t)=\cos(t\mathbf{L}_{\diamond})\varphi(u,0)$, assuming $\left.\frac{\partial\varphi(u,t)}{\partial t}\right|_{t=0}=0$. The following theorem establishes that, for Cartesian PGs, the error bound on the stability of our model admits the integrated sum of stability over the factor graphs.
Theorem 2.6.
[…]
$\|\varphi(u,t)-\tilde{\varphi}(u,t)\|=\sum_{p=1}^{P}{\mathcal{O}(\varepsilon_{p})}.$ (10)
Theorem 2.6 states that the solution of the So-TPDEG (3) remains stable with respect to the magnitude of perturbations in the factor graphs. This robustness is particularly important when the factor graphs are subject to errors.
2.3 Over-smoothing Analysis
Over-smoothing in GNNs is commonly characterized by the decay of Dirichlet energy as the number of layers increases. This phenomenon has been theoretically studied in prior works [24, 25, 15] and can be formulated for the output of a continuous GNN [15]. Specifically, consider the GNN output $\mathbf{U}_{t}=[\mathbf{u}(t)_{1},\ldots,\mathbf{u}(t)_{N}]^{\top}$, where $\mathbf{u}(t)_{i}$ represents the feature vector of node $i$, which has degree $\text{deg}_{i}$. Inspired by the previous studies [10], the Dirichlet energy can then be defined as $\lim_{t\rightarrow\infty}{E(\mathbf{U}_{t})}\rightarrow 0$, where
$\begin{split}E(\mathbf{U})\vcentcolon=\frac{1}{2}\sum_{(i,j)\in\mathcal{E}}{\left\|\frac{\mathbf{u}_{i}}{\sqrt{\text{deg}_{i}}}-\frac{\mathbf{u}_{j}}{\sqrt{\text{deg}_{j}}}\right\|^{2}}=\operatorname{\operatorname{tr}}(\mathbf{U}^{\top}\hat{\mathbf{L}}\mathbf{U}).\end{split}$ (11)
Motivated by (11), we generalize the concept of Dirichlet energy to tensors as follows:
Definition 2.7.
(Tensorial Dirichlet Energy [10]). Let $\{\hat{\mathbf{L}}_{p}\}_{p=1}^{P}$ denote the normalized factor Laplacians. For a tensor $\underline{\mathbf{U}}\in\mathbb{R}^{N_{1}\times\cdots\times N_{P}\times F}$, the Tensorial Dirichlet Energy is defined by:
$E(\underline{\mathbf{U}})\coloneqq\frac{1}{P}\sum_{f=1}^{F}\sum_{p=1}^{P}\mathrm{tr}\!\Big{(}\tilde{\mathbf{U}}_{f_{(p)}}^{\top}\hat{\mathbf{L}}_{p}\tilde{\mathbf{U}}_{f_{(p)}}\Big{)},$ (12)
where $\tilde{\mathbf{U}}_{f_{(p)}}=\underline{\mathbf{U}}[:,\ldots,:,f]_{(p)}\in\mathbb{R}^{N_{p}\times\prod_{i\neq p}^{P}N_{i}}$ represents the $p$-th mode matricization of the $f$-th slice of $\underline{\mathbf{U}}$ along the $(P+1)$-th dimension.
Drawing on prior work in the GSP [26, 27, 28], the tensorial Dirichlet energy $E(\underline{\mathbf{U}})$ defined in (12) can alternatively be expressed in matrix form as $E(\underline{\mathbf{U}})=\mathrm{tr}\!\Big{(}\tilde{\mathbf{U}}^{\top}\hat{\mathbf{L}}\tilde{\mathbf{U}}\Big{)}$, where $\tilde{\mathbf{U}}\in\mathbb{R}^{\left[\prod_{p=1}^{P}N_{p}\right]\times F}$, with columns given by $\tilde{\mathbf{U}}[:,f]=\operatorname{\operatorname{vec}}(\underline{\mathbf{U}}[:,\ldots,:,f])$, and where the normalized PG Laplacian is defined as $\hat{\mathbf{L}}\coloneqq\frac{1}{P}\oplus_{p=1}^{P}\hat{\mathbf{L}}_{p}=\oplus_{p=1}^{P}\frac{\hat{\mathbf{L}}_{p}}{P}$, implying a useful spectral property that the spectrum of $\hat{\mathbf{L}}$ spans the interval of $[0,2]$ [10]. From (12), we note that the separability of Cartesian heat kernels naturally extends to analyzing over-smoothing across the different modes of the data. This property is particularly advantageous when the factor graphs exhibit diverse characteristics. We now proceed with a formal analysis of over-smoothing in our model. Specifically, for the $l$-th layer, consisting of $H_{l}$ hidden MLP layers, nonlinear activations $\sigma(\cdot)$, and learnable weight matrices $\{\mathbf{W}_{lh}\}_{h=1}^{H_{l}}$, we define:
$\begin{split}&\mathbf{X}_{l+1}\vcentcolon=f_{l}(\mathbf{X}_{l}),\>\>\>f_{l}(\mathbf{X})\vcentcolon=\text{MLP}_{l}(\cos(\hat{\mathbf{L}}^{(t)})\mathbf{X}),\\ &\text{MLP}_{l}(\mathbf{X})\vcentcolon=\sigma(\ldots\sigma(\sigma(\mathbf{X})\mathbf{W}_{l1})\mathbf{W}_{l2}\ldots\mathbf{W}_{lH_{l}}),\end{split}$ (13)
Here, $\mathbf{X}_{0}$ denotes the initial node feature matrix, and $f_{l}(\cdot)$ represents a generalized form of (5). The term $cos(\hat{\mathbf{L}}^{(t)})$ is defined as $\hat{\mathbf{L}}:=\frac{1}{P}\oplus_{p=1}^{P}{\hat{\mathbf{L}}_{p}}$, where $t^{(p)}$ corresponds to the receptive field of the $p$-th factor graph. The following theorem then provides the formal criteria for over-smoothing.
Theorem 2.8.
For the product Laplacian $\hat{\mathbf{L}}:=\frac{1}{P}\oplus_{p=1}^{P}{\hat{\mathbf{L}}_{p}}$ with domain-specific receptive field $t^{(i)}$ for the $i$-th factor graph and the activations $\sigma(\cdot)$ in (13) being ReLU or Leaky ReLU, we have:
[…]
where
$\cos{(t\lambda^{\diamond}_{\phi})}=\sum_{\begin{subarray}{c}S\subseteq\{1,\dots,P\}\\ |S|\ \mathrm{even}\end{subarray}}(-1)^{|S|/2}\prod_{p\in S}\sin(t\lambda^{(p)}_{i_{p}})\prod_{p\notin S}\cos(t\lambda^{(p)}_{i_{p}})$. (15)
Here, $s\vcentcolon=\sup_{l\in\mathbb{N}_{+}}{s_{l}}$ and $s_{l}\vcentcolon=\prod_{h=1}^{H_{l}}{s_{lh}}$ with $s_{lh}$ being the square of maximum singular value of $\mathbf{W}^{\top}_{lh}$, and $\lambda^{(p)}_{i_{p}}$ is the $i_{p}$-th eigenvalue of the $p$-th normalized factor graph such that $\cos^{2}{(t\lambda^{\diamond}_{\phi})}=\max_{i}{\cos^{2}{(t\lambda^{\diamond}_{i})}}$, where $\lambda^{\diamond}_{i}$ is the $i$-th non-zero eigenvalue of $\hat{\mathbf{L}}_{\diamond}$ and $\lambda^{\diamond}_{\phi}=\sum_{p=1}^{P}{\lambda^{(p)}_{i_{p}}}$. Besides, $\cos^{2}{(t\lambda^{\diamond}_{\phi})}=\max_{i}{\cos^{2}{(t\lambda^{\diamond}_{i})}}$, where $\lambda^{\diamond}_{i}$ is the $i$-th non-zero eigenvalue of $\hat{\mathbf{L}}_{\diamond}$.
Corollary 2.9.
If $s\cos^{2}{(t\lambda^{\diamond}_{\phi})}<1$ and $l\rightarrow\infty$, $E(\mathbf{X}_{l})$ exponentially converges to 0, leading to:
$t\in\bigcup_{k\in\mathbb{Z}}\frac{1}{2\lambda^{\diamond}_{\phi}}(2k\pi+\arccos(\frac{2}{s}-1),2(k+1)\pi-\arccos(\frac{2}{s}-1)).$ (16)
Theorem 2.8 and Corollary 2.9 demonstrate that the oversmoothing rate is influenced by three key factors: the learnable weight matrices (through $s$), the spectral characteristics of the factor and product graph (via $\lambda^{\diamond}_{\phi}$), and the graph receptive field $t$. Unlike first-order TPDEGs [10], which yield only a single lower bound on $t$ and consequently suppress high-frequency components, (16) produces infinitely valid intervals. This flexibility enables control over oversmoothing by tuning $t$ and naturally leads to oscillatory solutions acting as an inherent band-pass filter on the factor graphs.
3 Conclusion and Discussion
In this work, we introduced So-TPDEG, a second-order tensorial PDE framework designed to process data over multiple interacting (product) graphs, supported by a rigorous theoretical foundation. Leveraging this formulation, we develop a second-order continuous product GNN, a novel model that effectively solves these PDEs while providing a principled treatment of stability and over-smoothing. Compared to current first-order TPDEG models, So-TPDEG does not dampen the medium to high-frequency components (because it is not a low-pass graph filter [8]), which might be applicable in some real-world applications with periodic patterns or fine-grained details [19, 20, 21].